A three-month (90 day) Treasury bill with a $10,000 face value is quoted at a discount...

  1. A three-month (90 day) Treasury bill with a $10,000 face value is quoted at a discount of 0.60.
    1. What is the implied price of the Treasury bill?
    2. What is the implied rate of return that an investor can earn over the 90-day period?
Answer #1
a) Implied price = 10000*(1-0.60%) = $           9,940
b) Implied rate of return = 10000/9940-1 = 0.6036%